OpenStudy (alibaby):

What is the sum of 4.2 × 10^5 and 5.3 × 10^5? 95 × 10^5 95 × 10^10 9.5 × 10^5 9.5 × 10^10

OpenStudy (alibaby):

@nermaljean99

OpenStudy (nermaljean99):

alrighty. break it down. 4.2*5.3

OpenStudy (alibaby):

22.26

OpenStudy (nermaljean99):

Yep! Now for the next part.

OpenStudy (nermaljean99):

10^5*10^5

OpenStudy (alibaby):

10000000000

OpenStudy (nermaljean99):

Haha is that an answer choice?

OpenStudy (alibaby):

B?

OpenStudy (danjs):

It says Sum, add them, not multiply

OpenStudy (3mar):

+ not *

OpenStudy (nermaljean99):

Yes. gosh. that stupid mistake...

OpenStudy (alibaby):

so B

OpenStudy (nermaljean99):

yes sorry about the first part. I have dyslexia

OpenStudy (3mar):

So it will be\[(4.2\times10^5) + (5.3\times10^5)=10^5(4.2+5.3)=9.5\times10^5\] The third choice. Hope that helps

OpenStudy (3mar):

No problem.